\begin{thebibliography}{99}

\bibitem{BHS} Andreas R. Blass, Jeffry L. Hirst, and Stephen G. Simpson,
Logical analysis of some theorems of combinatorics and topological
dynamics, in Stephen G. Simpson, ed., Logic and Combinatorics,
Contemporary Mathematics 65, American Mathematical Society,
Providence, RI, 1987, 125--156.

\bibitem{Car} Lorenzo Carlucci, Restrictions on Hindman's Theorem: an
overview, in Liesbeth De Mol, Andreas Weiermann, Florin Manea, and
David Fern\'andez-Duque, eds., Connecting with Computability: CiE
2021, Lecture Notes in Computer Science 12813, Springer, Cham, 2021,
94--105.

\bibitem{CKLZ} Lorenzo Carlucci, Leszek A. Ko\l{}odzieczyk, Francesco
Lepore, and Konrad Zdanowski, New bounds on the strength of some
restrictions of Hindman's Theorem, Computability 9 (2020), no.\ 2,
139--153. 

\bibitem{CDHJSW} Barbara F. Csima, Damir D. Dzhafarov, Denis
R. Hirschfeldt, Carl G. Jockusch, Jr., Reed Solomon, and Linda Brown
Westrick, The reverse mathematics of Hindman's Theorem for sums of
exactly two elements, Computability 8 (2019), no.\ 3-4, 253--263.

\bibitem{DJSW} Damir D. Dzhafarov, Carl G. Jockusch, Jr., Reed
Solomon, and Linda Brown Westrick, Effectiveness of Hindman's
Theorem for bounded sums, in Adam Day, Michael Fellows, Noam
Greenberg, Bakhadyr Khoussainov, Alexander Melnikov, and Frances
Rosamond, eds., Computability and Complexity: Essays Dedicated to
Rodney G. Downey on the Occasion of His 60th Birthday, Lecture Notes
in Computer Science 10010, Springer, Cham, 2017, 134--142.

\bibitem{Hindman} Neil Hindman, Finite sums from sequences within
cells of a partition of $N$, Journal of Combinatorial Theory Series A
17 (1974), no.\ 1, 1--11.

\bibitem{HLS} Neil Hindman, Imre Leader, and Dona Strauss, Open
problems in partition regularity, Combinatorics, Probability and
Computing 12 (2003), no.\ 5-6, 571--583.

\bibitem{Hbook} Denis R. Hirschfeldt, Slicing the Truth, Lecture Notes
Series, Institute for Mathematical Sciences, National University of
Singapore 28, World Scientific Publishing Co. Pte. Ltd., Hackensack,
NJ, 2015.

\bibitem{J} Carl G. Jockusch, Jr., Ramsey's theorem and recursion
theory, The Journal of Symbolic Logic 37 (1972), no.\ 2, 268--280.

\bibitem{Jfpf} Carl G. Jockusch, Jr., Degrees of functions with no
fixed points, in Jens Erik Fenstad, Ivan T. Frolov and Risto Hilpinen,
eds., Logic, Methodology, and Philosophy of Science VIII
(Moscow, 1987), Studies in Logic and the Foundations of Mathematics 126,
North-Holland, Amsterdam, 1989, 191--201.

\bibitem{L1} Jiayi Liu, RT$^2_2$ does not imply WKL$_0$, The Journal
of Symbolic Logic 77 (2012), no.\ 2, 609--620.

\bibitem{L2} Lu Liu, Cone avoiding closed sets, Transactions of the
American Mathematical Society 367 (2015), no.\ 3, 1609--1630.

\bibitem{P} Ludovic Patey, Somewhere over the rainbow Ramsey theorem
for pairs, arXiv:1501.07424.

\bibitem{P2} Ludovic Patey, On combinatorial weaknesses of Ramseyan
principles,
https://ludovicpatey.com/media/research/combinatorial-weaknesses-draft.pdf.

\bibitem{R} Andrei Rumyantsev, Infinite computable version of Lov\'asz
Local Lemma, arXiv:1012.0557.

\bibitem{RS} Andrei Rumyantsev and Alexander Shen, Probabilistic
constructions of computable objects and a computable version of
Lov\'asz Local Lemma, Fundamenta Informaticae 132 (2014), no.\ 1,
1--14.

\bibitem{SS} David Seetapun and Theodore A. Slaman, On the strength of
Ramsey's Theorem, Notre Dame Journal of Formal Logic 36 (1995),
no.\ 4, 570--582.

\bibitem{Sbook} Stephen G. Simpson, Subsystems of Second Order
Arithmetic, First edition, Perspectives in Mathematical Logic,
Springer-Verlag, Berlin, 1999.

\bibitem{W} Wei Wang, Some logically weak Ramseyan theorems,
Advances in Mathematics 261 (2014) 1--25.

\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{amsmath, amsthm, amssymb, latexsym, enumerate}

\usepackage[all]{xy}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}[thm]{Lemma}
\newtheorem{cor}[thm]{Corollary}

\theoremstyle{definition}

\newtheorem{defn}[thm]{Definition}
\newtheorem{q}[thm]{Question}

\renewcommand{\leq}{\leqslant}
\renewcommand{\geq}{\geqslant}

\DeclareMathOperator{\sg}{\textit{sg}}

\DeclareMathOperator{\vsg}{\textit{vsg}}

\DeclareMathOperator{\fs}{\textit{fs}}

\title{Thin Set Versions of Hindman's Theorem}

\renewcommand{\footnotemark}{}

\author{Denis R. Hirschfeldt and Sarah C. Reitzes\\Department of
Mathematics, University of Chicago\thanks{The authors were partially
supported by Focused Research Group grant DMS-1854279 from the
National Science Foundation of the United States. Hirschfeldt was also
partially support by NSF grant DMS-1600543, and Reitzes by NSF grant
DGE-1746045.}}

\begin{document}

\maketitle

\section{Introduction}

This paper is part of a line of research on the
computability-theoretic and reverse-mathematical strength of versions
of Hindman's Theorem~\cite{Hindman} that began with the work of Blass,
Hirst, and Simpson~\cite{BHS}, and has seen considerable interest
recently. We assume basic familiarity with computability theory and
reverse mathematics, at the level of the background material in
\cite{Hbook}, for instance. On the reverse mathematics side, the two
major systems with which we will be concerned are RCA$_0$, the usual
weak base system for reverse mathematics, which corresponds roughly to
computable mathematics; and ACA$_0$, which corresponds roughly to
arithmetic mathematics. For principles $P$ of the form $(\forall
X)\, [\Phi(X) \rightarrow (\exists Y)\, \Psi(X,Y)]$, we call any $X$
such that $\Phi(X)$ holds an \emph{instance} of $P$, and any $Y$ such
that $\Psi(X,Y)$ holds a \emph{solution} to $X$.

We begin by introducing some related combinatorial principles.  For a
set $S$, let $[S]^n$ be the set of $n$-element subsets of
$S$. \emph{Ramsey's Theorem} (RT) is the statement that for every $n$
and every coloring of $[\mathbb N]^n$ with finitely many colors, there
is an infinite set $H$ that is \emph{homogeneous} for $c$, which means
that all elements of $[H]^n$ have the same color. There has been a
great deal of work on computability-theoretic and reverse-mathematical
aspects of versions of Ramsey's Theorem, such as RT$^n_k$, which is RT
restricted to colorings of $[\mathbb N]^n$ with $k$ many colors. (See
e.g.\ \cite{Hbook}.)

The Thin Set Theorem is another variant of Ramsey's Theorem that has
been studied from this perspective. It follows easily from Ramsey's
Theorem itself.

\begin{defn}
\emph{Thin Set Theorem} (TS): For every $n$ and every coloring $c :
[\mathbb  N]^n \rightarrow \mathbb N$, there is an infinite set $T
\subseteq \mathbb N$ and an $i$ such that $c(s) \neq i$ for all $s \in
[T]^n$. We call such a set $T$ a \emph{thin set} for $c$. TS$^n$ is
the restriction of TS to colorings of $[\mathbb N]^n$.
\end{defn}

Jockusch~\cite{J} showed that there is a computable instance of
RT$^3_2$ such that any solution computes the halting problem
$\emptyset'$. As shown by Simpson~\cite{Sbook}, Jockusch's
construction can also be used to prove that RT$^3_2$ (and hence RT)
implies ACA$_0$ over RCA$_0$. Wang~\cite{W} showed that TS, on the
other hand, does not have this much power. Indeed, it has a property
known as strong cone avoidance, which implies in particular that for
every coloring $c : [\mathbb N]^n \rightarrow \mathbb N$ and every
noncomputable $X$, there is an infinite thin set for $c$ that does not
compute $X$. It also follows from strong cone avoidance that TS does
not imply ACA$_0$ over RCA$_0$.

As shown by Seetapun~\cite{SS}, RT$^2_k$ also fails to imply
ACA$_0$. Indeed, Liu~\cite{L1,L2} showed that it does not imply the
weaker system WKL$_0$, which consists of RCA$_0$ together with Weak
K\"onig's Lemma, or the even weaker system WWKL$_0$ consisting of
RCA$_0$ together with Weak Weak K\"onig's Lemma. Patey~\cite{P2}
showed that the same is true of TS.

We now turn to Hindman's Theorem. For a set $S \subseteq \mathbb N$,
let $\fs(S)$ be the set of sums of nonempty finite sets of distinct
elements of $S$.

\begin{defn}
\emph{Hindman's Theorem} (HT): For every coloring of $\mathbb N$ with
finitely many colors, there is an infinite set $S \subseteq \mathbb N$
such that all elements of $\fs(S)$ have the same color.
\end{defn}

Blass, Hirst, and Simpson~\cite{BHS} showed that such an $S$ can
always be computed in the $(\omega+1)$st jump of the coloring, and
that there is a computable coloring such that every such $S$
computes $\emptyset'$. By analyzing these proofs they showed that HT
is provable in ACA$_0^+$ (the system consisting of RCA$_0$ together
with the statement that $\omega$th jumps exist) and implies ACA$_0$
over RCA$_0$. The exact computability-theoretic and
reverse-mathematical strength of HT remains open.

There has recently been interest in studying restricted versions of HT
such as the following. (See e.g.\ \cite{Car}.)

\begin{defn}
HT$^{\leq n}$ is HT restricted to sums of at most $n$ many elements,
and HT$^{=n}$ is HT restricted to sums of exactly $n$ many
elements. HT$^{\leq n}_k$ and HT$^{=n}_k$ are the corresponding
restrictions to colorings with $k$ many colors.
\end{defn}

Dzhafarov, Jockusch, Solomon, and Westrick~\cite{DJSW} showed that
HT$^{\leq 3}_3$ implies ACA$_0$ over RCA$_0$. Carlucci,
Ko\l{}odzieczyk, Lepore, and Zdanowski~\cite{CKLZ} did the same for
HT$^{\leq 2}_4$. These principles are also complex in a more heuristic
sense: There is no known way to prove even HT$^{\leq 2}_2$ other than
to give a proof of the full HT, which has led Hindman, Leader, and
Strauss~\cite{HLS} to ask whether every proof of HT$^{\leq 2}$ is also
a proof of HT. This question can be formalized by asking whether
HT$^{\leq 2}$ (or HT$^{\leq 2}_2$) implies HT, say over RCA$_0$. A
related open question is whether HT$^{\leq 2}_2$ is provable in
ACA$_0$.

The principle HT$^{=2}$ is quite different, as HT$^{=2}_k$ follows
easily from RT$^2_k$. Indeed, it was not clear even whether this
principle is computably true until the work of Csima, Dzhafarov,
Hirschfeldt, Jockusch, Solomon, and Westrick~\cite{CDHJSW}, who showed
that it is not, and that indeed there is a computable instance of
HT$^{=2}_2$ with no $\Sigma^0_2$ solutions. (The same had been shown
for RT$^2_2$ by Jockusch~\cite{J}, who also showed that every
computable instance of RT$^2_2$ has a $\Pi^0_2$ solution, which
implies that the same is true of HT$^{=2}_2$.) They also showed that
there is  a computable instance of HT$^{=2}_2$ such that every
solution has DNC degree relative to $\emptyset'$, and adapted this
proof to show that HT$^{=2}_2$ implies the principle RRT$^2_2$, a
version of the Rainbow Ramsey Theorem, over RCA$_0$. (See Section
\ref{hard} for definitions.)

In this paper, we study further versions of Hindman's Theorem,
obtained by combining HT and its variants with the Thin Set Theorem.

\begin{defn}
thin-HT: For every coloring $c : \mathbb N \rightarrow \mathbb N$,
there is an infinite set $S \subseteq \mathbb N$ such that $\fs(S)$ is
thin for $c$. We definite restrictions such as thin-HT$^{\leq n}$
analogously.
\end{defn}

In Section~\ref{encoding}, we give similar lower bounds on the
complexity of thin-HT as Blass, Hirst, and Simpson~\cite{BHS} gave for
HT, which suggests that thin-HT behaves like HT at least to some
extent. Indeed, it seems possible that thin-HT is equivalent to HT
over RCA$_0$. The situation for restricted versions is different,
however. Clearly, thin-HT$^{=n}$ follows from TS$^n$, but in fact so
does thin-HT$^{\leq n}$, due to the following fact.

\begin{lem}
For each $n$ and $k$, the following holds in $\textup{RCA}_0 +
\textup{TS}^n$: Given $c_i : [\mathbb N]^{m_i} \rightarrow \mathbb N$
for $i \leq k$, with $m_i \leq n$ for all $i \leq k$, there is a
single infinite set $T$ and a $j$ such that $c_i(s) \neq j$ for each
$c_i$ and each $s \in [T]^{m_i}$ with $i \leq k$.
\end{lem}

\begin{proof}
We use the fact that TS$^n$ implies TS$^m$ for each $m<n$, and proceed
by external induction to prove the stronger assertion that for each $j
\leq k$, $\textup{RCA}_0 + \textup{TS}^n$ proves that there is an
infinite set $T$ and an infinite set $C$ such that $c_i(s) \notin C$
for each $c_i$ and each $s \in [T]^{m_i}$ with $i \leq j$.

We do the base and inductive cases simultaneously. For $j+1>0$, assume
that that the assertion holds for $j$ and let $T$ and $C$ be as
above. For $j+1=0$, let $T=C=\mathbb N$. Define $d : [T]^{m_{j+1}}
\rightarrow \mathbb N$ as follows. Partition $C$ into infinitely many
infinite sets $A_0,A_1,\dots$. Let $d(s) = 0$ if either $c_{j+1}(s)
\in A_0$ or $c_{j+1}(s) \notin C$, and for $i>0$, let $d(s)=i$ if
$c_{j+1}(s) \in A_i$. By TS$^{m_{j+1}}$, there is an infinite $U
\subseteq T$ that is thin for $d$. Let $i \notin d([U]^{m_{j+1}})$ and
let $D=A_i$. Then $U$ and $D$ are infinite sets such that $c_i(s)
\notin D$ for each $c_i$ and each $s \in [U]^{m_i}$ with $i \leq j+1$.
\end{proof}

This lemma allows us to get thin-HT$^{\leq n}$ from TS$^n$ by taking a
coloring $c: \mathbb N \rightarrow \mathbb N$ and considering the
colorings that map $\{a_0,\ldots,a_j\}$ to $c(a_0+\cdots+a_j)$ for
each $j<n$.

There are also differences that have nothing to do with computability
theory and reverse mathematics between thin-HT$^{\leqslant n}$ on the
one hand, and thin-HT and HT$^{\leqslant n}$ on the other. The former
remains true if we allow sums of non-distinct elements, but it is not
difficult to show that the latter two do not. Similarly, the former
remains true for colorings $S \rightarrow \mathbb N$, where $S
\subseteq \mathbb N$ is any infinite set, while the latter two again
do not.

Nevertheless, even thin-HT$^{=2}$ still has a significant level of
complexity. In Section~\ref{hard}, we show that all of the lower
bounds mentioned above obtained in \cite{CDHJSW} for HT$^{=2}$ still
hold for thin-HT$^{=2}$.

In Section~\ref{open} we mention some open questions arising from our
results, and briefly discuss version of HT obtained by combining it
with thin set theorems for colorings with finitely many colors.

\section{Encoding $\emptyset'$ into thin-HT}
\label{encoding}

In this section, we show how to build on the proof of Theorem 2.2 of
Blass, Hirst, and Simpson~\cite{BHS}, which shows that there is a
computable instance of HT such that every solution computes
$\emptyset'$, to show that the same is true of thin-HT. We then derive
a reverse-mathematical consequence of our proof.

\begin{thm}
There is a computable instance of \textup{thin-HT} such that every
solution computes $\emptyset'$.
\end{thm}

\begin{proof}
As in the proof of Theorem 2.2 of \cite{BHS}, we write each number
$x>0$ as $2^{n_0}+\cdots+2^{n_k}$ with $n_0<\cdots<n_k$, and define
$\lambda(x)=n_0$ and $\mu(x)=n_k$. A set $S$ \emph{has $2$-apartness}
if for every $x,y \in S$ with $x<y$, we have
$\mu(x)<\lambda(y)$. Lemma 4.1 of \cite{BHS} shows that from any
infinite $S$ we can compute an infinite set $T$ with $2$-apartness
such that $\fs(T) \subseteq \fs(S)$ (and hence if $\fs(S)$ is thin for
a coloring, so is $\fs(T)$).

Let $x=2^{n_0}+\cdots+2^{n_k}$ with $n_0<\cdots<n_k$. Say that
$(n_i,n_{i+1})$ is a \emph{short gap} in $x$ if there is an $m<n_i$
such that $m \notin \emptyset'[n_{i+1}]$ but $m \in \emptyset'$. Say
that $(n_i,n_{i+1})$ is a \emph{very short gap} in $x$ if there is an
$m<n_i$ such that $m \notin \emptyset'[n_{i+1}]$ but $m \in
\emptyset'[n_k]$. Let $\sg(x)$ and $\vsg(x)$ be the numbers of short
gaps and very short gaps in $x$, respectively. Note that $\sg$ is not
a computable function, but $\vsg$ is.

Fix a bijection between $\mathbb N$ and the set of pairs $(p,i)$ where
$p$ is prime and $1 \leq i < p$, and identify $\mathbb N$ with this
set via this bijection. Define the coloring $c$ by letting
$c(x)=(p,i)$ where $p$ is the least prime that does not divide
$\vsg(x)$ and $\vsg(x) = i \bmod p$. We say that $x$ \emph{has color
$(p,i)$} if $c(x)=(p,i)$, and we also say that $x$ \emph{has color}
$(p,0)$ or $(p,p)$ if it has color $(q,i)$ for some $q>p$, i.e., if
every prime less than or equal to $p$ divides $\vsg(x)$.

Let $Y$ be such that $\fs(Y)$ is an infinite thin set for $c$. We can
assume that $Y$ has $2$-apartness, by Lemma 4.1 of \cite{BHS}, as
mentioned above. This condition ensures that if $x,y \in \fs(Y)$ and
$\mu(x)<\lambda(y)$, and we express $x$ and $y$ as sums of sets $F$
and $G$ of distinct elements of $Y$, respectively, then $F$ and $G$
are disjoint, and hence $x+y \in \fs(Y)$.  Say that $S \subseteq
\fs(Y)$ is \emph{$\lambda$-bounded} if there is a bound on the values
of $\lambda(x)$ for $x \in S$ (which includes the case
$S=\emptyset$). Note that $\fs(Y)$ itself is not
$\lambda$-bounded. Note also that the union of finitely many
$\lambda$-bounded sets is $\lambda$-bounded. Say that a color $j$ is
\emph{almost absent} from $\fs(Y)$ if the set of $x \in \fs(Y)$ that
have color $j$ is $\lambda$-bounded. (This definition includes the
case $j=(p,0)$, or equivalently $j=(p,p)$.)

\begin{lem}
There are $p$ and $0 \leq i < p$ such that $(p,i+1)$ is almost absent
from $\fs(Y)$ but $(p,i)$ is not.
\end{lem}

\begin{proof}
Let $p$ be least such that there is a $j$ for which $(p,j)$ is almost
absent from $\fs(Y)$, which exists since $\fs(Y)$ is thin. If $p=2$
then $(p,j+1)$ cannot be almost absent, since every number has color
$(p,j)$ or $(p,j+1)$. Now suppose that $p>2$ and $q$ is the preceding
prime. Since $(q,0)$ is not almost absent from $\fs(Y)$ and every
number that has color $(q,0)$ has color $(p,j)$ for some $j$, there
is some $k$ such that $(p,k)$ is not almost absent. In either case,
since having color $(p,0)$ is the same as having color $(p,p)$, the
lemma follows.
\end{proof}

Fix $p$ and $i$ as in the above lemma.

\begin{lem}
Let $1 \leq j <p$. Then $S=\{x \in \fs(Y) : \sg(x) = j \bmod p\}$ is
$\lambda$-bounded.
\end{lem}

\begin{proof}
Suppose $S$ is not $\lambda$-bounded. Let $q_0<\cdots<q_{m-1}$ be the
primes less than $p$. Since there are only finitely many sequences
$(k_0,\ldots,k_{m-1})$ with $k_i < q_i$, there is such a sequence for
which $T=\{x \in S : (\forall \ell<m)\, \sg(x) = k_\ell \bmod
q_\ell\}$ is not $\lambda$-bounded.

Since $j \neq 0 \bmod p$, and hence $q_0 \cdots q_{m-1} j \neq 0 \bmod
p$, there is a multiple $n$ of $q_0 \cdots q_{m-1}$ such that $nj = 1
\bmod p$ (where $q_0 \cdots q_{m-1} = 1$ if $p=2$). Since $T$ is not
$\lambda$-bounded, there are $x_0<\cdots<x_{n-1} \in T$ such that each
$\lambda(x_{k+1})$ is sufficiently large relative to $\mu(x_k)$ to
ensure that $(\mu(x_k),\lambda(x_{k+1}))$ is not a short gap. Then the
short gaps in $x_0+\cdots+x_{n-1}$ are exactly the short gaps in
$x_0,\ldots,x_{n-1}$, so
$\sg(x_0+\cdots+x_{n-1})=\sg(x_0)+\cdots+\sg(x_{n-1})$. The latter is
equal to $nj \bmod p = 1 \bmod p$, since each $x_\ell$ is in $S$, and
is also equal to $nk_\ell \bmod q_\ell$ for each $\ell<m$, and hence
equal to $0 \bmod q_\ell$ for each $\ell<m$, since $n = 0 \bmod
q_\ell$.

Since $(p,i)$ is not almost absent from $\fs(Y)$, there is a $y \in
\fs(Y)$ that has color $(p,i)$ such that $\lambda(y) > \mu(x_{n-1})$,
and every number less than $\mu(x_{n-1})$ that is in $\emptyset'$ is
already in $\emptyset'[\lambda(y)]$. Note that $\vsg(y) = 0 \bmod
q_\ell$ for each $\ell<m$, as otherwise $c(y)$ would be of the form
$(q_\ell,k)$ for some $1 \leq k < q_\ell$. Now $\vsg(x_0 + \cdots +
x_{n-1} + y) = \vsg(y) + \sg(x_0 + \cdots + x_{n-1})$, which is equal
to $i+1 \bmod p$, and to $0 \bmod q_\ell$ for all $\ell<m$. So $x_0 +
\cdots + x_{n-1} + y$ has color $(p,i+1)$. As we can choose $x_0$ so
that $\lambda(x_0)$ is arbitrarily large, $(p,i+1)$ is not almost
absent from $\fs(Y)$, contradicting the choice of $i$.
\end{proof}

So by removing finitely many elements from $Y$ if needed, we can
assume that $p$ divides $\sg(x)$ for all $x \in \fs(Y)$. We can now
argue as in the proof of Claim 2 in the proof Theorem 2.2 of
\cite{BHS} to compute $\emptyset'$ from $Y$: Given $n$, find $x,y \in
Y$ such that $x<y$ and $n < \mu(x)$. The short gaps in $x+y$ are the
ones in $x$, the ones in $y$, and possibly $(\mu(x),\lambda(y))$. But
if the latter is a short gap, then $\sg(x+y)=\sg(x)+\sg(y)+1$, which
is impossible since $p$ divides all three numbers. Thus $n \in
\emptyset'$ if{}f $n \in \emptyset'[\lambda(y)]$.
\end{proof}

The above proof can be carried out in relativized form in RCA$_0$
except for two issues: One is that in RCA$_0$ we cannot show that the
union of finitely many $\lambda$-bounded sets is $\lambda$-bounded,
which in general requires the $\Pi^0_1$-bounding principle. Another is
that being almost absent is a $\Sigma^0_2$ condition, so we cannot
conclude in RCA$_0$ that there is a least $p$ such that there is a $j$
for which $(p,j)$ is almost absent from $\fs(Y)$. Since
$\Pi^0_1$-bounding follows from $\Sigma^0_2$-induction over RCA$_0$,
adding the latter to RCA$_0$ is sufficient to get around these issues,
so we have the following.

\begin{thm}
\textup{thin-HT} implies \textup{ACA}$_0$ over $\textup{RCA}_0 +
\textup{I}\Sigma^0_2$.
\end{thm}

We do not know whether the use of I$\Sigma^0_2$ in this theorem can be
removed.

\section{Hard Instances of thin-HT$^{=2}$}
\label{hard}

In this section, we show that all the lower bounds on the complexity
of HT$^{=2}_2$ obtained by Csima, Dzhafarov, Hirschfeldt, Jockusch,
Solomon, and Westrick~\cite{CDHJSW} still hold for thin-HT$^{=2}$. (Of
course, all upper bounds on the complexity of HT$^{=2}_2$
automatically hold for thin-HT$^{=2}$, as the latter follows easily
from the former.) As in that paper, we use the computable version of
the Lov{\'a}sz Local Lemma due to Rumyantsev and Shen~\cite{R,RS}. In
particular, we use the following consequence of Corollary 7.2 in
\cite{RS} given in \cite{CDHJSW}, with an addendum on uniformity as
noted at the end of Section 4 of \cite{CDHJSW}. This uniformity, which
in \cite{CDHJSW} is used only to obtain results on Weihrauch
reducibility, will be essential in all our results, as their proofs
will require applying Theorem \ref{rsthm} infinitely often.

\begin{thm}[essentially Rumyantsev and Shen~\cite{RS}] 
\label{rsthm}
For each $q \in (0,1)$ there is an $M$ such that the following
holds. Let $F_0,F_1,\ldots$ be a computable sequence of finite sets,
each of size at least $M$. Suppose that for each $m \geq M$ and $n$,
there are at most $2^{qm}$ many $j$ such that $|F_j|=m$ and $n \in
F_j$, and that there is a computable procedure $P$ for determining the
set of all such $j$ given $m$ and $n$. Then there is a computable $c :
\mathbb N \rightarrow 2$ such that for each $j$ the set $F_j$ is not
homogeneous for $c$. Furthermore, $c$ can be obtained uniformly
computably from $F_0,F_1,\ldots$ and $P$ (for a fixed $q$).
\end{thm}

We will also rely in this section on arguments in \cite{CDHJSW} when
they carry through in this case in an entirely analogous way.

We now introduce a notion of largeness that will be key to our
iterated applications of Theorem \ref{rsthm}. As in \cite{CDHJSW}, we
will be diagonalizing against $\Sigma^0_2$ sets, so this notion will
be defined in terms of sets that are c.e.\ relative to $\emptyset'$.
For a set $A$ and a number $s$, we write $s+A$ for the set $\{s+a : a
\in A\}$. We write $W_e$ for the $e$th enumeration operator. Given $e$
and $s$, for each $x \in W^{\emptyset'}_e[s]$, let $t_x$ be the least
$t$ such that $x \in W^{\emptyset'}_e[u]$ for all $u \in
[t,s]$. (I.e., $t_x$ measures how long $x$ has been in
$W^{\emptyset'}_e$.) Order the elements of $W^{\emptyset'}_e[s]$ by
letting $x \prec y$ if either $t_x<t_y$ or both $t_x=t_y$ and
$x<y$. Let $E_e^n[s]$ be the set consisting of the least $n$ many
elements of $W^{\emptyset'}_e[s]$ under this ordering, or
$E_e^n[s]=[0,n)$ if $W^{\emptyset'}_e[s]$ has fewer than $n$ many
elements. If there is an $s$ such that $E_e^n[t] = E_e^n[s]$ for all
$t>s$ then let $E_e^n = E_e^n[s]$.

\begin{defn}
For a binary function $f$, say that a set $D$ is \emph{$f$-large} if
for all $e$ and $k$ such that $E_e^{f(e,k)}$ is defined, we have $|D
\cap (s+E_e^{f(e,k)})| \geq k$ for all sufficiently large $s$.
\end{defn}

Note that $\mathbb N$ is $g$-large for the function $g(e,k)=k$, and
that $f$-largeness is preserved under finite difference. The following
lemma captures the key property of this notion of largeness.

\begin{lem}
\label{mainlem}
From a binary function $f$ and an $f$-large set $D$, we can uniformly
compute a binary function $\widehat{f}$ and a splitting $D = D^0
\sqcup D^1$ such that each $D^i$ is $\widehat{f}$-large.
\end{lem}

Before proving this lemma, let us derive some of its consequences,
beginning with computability-theoretic lower bounds on the complexity
of thin-HT$^{=2}$. A function $f$ is \emph{diagonally noncomputable}
(\emph{DNC}) relative to an oracle $X$ if $f(e) \neq \Phi^X_e(e)$ for
all $e$ such that $\Phi^X_e(e)$ is defined, where $\Phi_e$ is the
$e$th Turing functional. A degree is \emph{DNC} relative to $X$ if it
computes a function that is DNC relative to $X$. An infinite set $A$
is \emph{effectively immune} relative to $X$ if there is an
$X$-computable function $f$ such that if $W^X_e \subseteq A$ then
$|W^X_e| < f(e)$.

\begin{thm}[Jockusch~\cite{Jfpf}]
\label{Jthm}
A degree is DNC relative to $X$ if and only if it computes a set that
is effectively immune relative to $X$.
\end{thm}

The proof of the following theorem shows how to obtain a hard
computable instance of thin-HT$^{=2}$ from Lemma \ref{mainlem}.

\begin{thm}
\label{dncthm}
There is a computable instance of \textup{thin-HT}$^{=2}$ such that
any solution is effectively immune relative to $\emptyset'$, and hence
has DNC degree relative to $\emptyset'$.
\end{thm}

\begin{proof}
Let $D_0=\mathbb N$ and $f_0(e,k)=k$. Given $D_n$ and $f_n$, let
$\widehat{f_n}$ and $D_n^i$ be as in Lemma \ref{mainlem}, let
$f_{n+1}=\widehat{f_n}$, and let $D_{n+1}=D_n^1$. Note that the $D_n$
are uniformly computable. Let $c(x)$ be the largest $n \leq x$ such
that $x \in D_n$. Then $c$ is a computable coloring of $\mathbb N$.
If $c(x)=n$ and $x>n$ then $x \in D_n$ but $x \notin D_m$ for $m>n$,
so $x \in D_n^0$. Thus for each $n$, we have that the difference
between $c^{-1}(n)$ and $D^0_n$ is finite, and hence $c^{-1}(n)$ is
$f_n$-large.

Let $S$ be a solution to $c$ as an instance of thin-HT$^{=2}$, and let
$n$ be such that $c(x+y) \neq n$ for all distinct $x,y \in S$. For any
$e$, if $|W_e^{\emptyset'}| \geq f_n(e,1)$ then $E_e^{f_n(e,1)}
\subseteq W_e^{\emptyset'}$ is defined, and hence $c^{-1}(n) \cap
(s+E_e^{f_n(e,1)}) \neq \emptyset$ for all sufficiently large $s$. In
other words, if $s$ is sufficiently large then there is an $x \in
E_e^{f_n(e,1)}$ such that $c(x+s)=n$. It follows that $E_e^{f_n(e,1)}
\nsubseteq S$, and hence $W_e^{\emptyset'} \nsubseteq S$, since
$E_e^{f_n(e,1)} \subseteq W_e^{\emptyset'}$. Thus we conclude that if
$W_e^{\emptyset'} \subseteq S$ then $|W_e^{\emptyset'}| <
f_n(e,1)$. Since $f_n(e,1)$ is computable as a function of $e$, it
follows that $S$ is effectively immune relative to $\emptyset'$, and
hence has DNC degree relative to $\emptyset'$.
\end{proof}

No infinite $\Sigma^0_2$ set can be effectively immune relative to
$\emptyset'$, so we have the following.

\begin{cor}
There is a computable instance of \textup{thin-HT}$^{=2}$ with no
$\Sigma^0_2$ solution.
\end{cor}

It follows that thin-HT is not provable in WKL$_0$, since the latter
has $\omega$-models consisting entirely of $\Delta^0_2$ sets. It was
noted in \cite{CDHJSW} that HT$^{=2}_2$ does not imply WKL$_0$, and
hence neither does thin-HT$^{=2}$. Thus thin-HT$^{=2}$ and WKL$_0$ are
incomparable over RCA$_0$. In fact, as mentioned in the introduction,
Patey~\cite{P2} showed that TS does not imply WKL$_0$, or even
WWKL$_0$, and we can easily adapt the proof of Theorem \ref{dncthm} to
thin-HT$^{=n}$ for any $n>2$, so we have the following.

\begin{cor}
For each $n>1$, both \textup{thin-HT}$^{=n}$ and
\textup{thin-HT}$^{\leq n}$ are incomparable with \textup{(W)WKL}$_0$
over \textup{RCA}$_0$.
\end{cor}

Arguing as in the proof of Corollary 3.6 of \cite{CDHJSW}, we have the
following.

\begin{cor}
There is a computable instance of \textup{thin-HT}$^{=2}$ such that
all solutions are hyperimmune.
\end{cor}

The reverse-mathematical analog of the existence of degrees that are
DNC over the jump is the principle 2-DNC, defined e.g.\ in Section 4
of \cite{CDHJSW}. Miller [unpublished] showed that 2-DNC is
equivalent, both over RCA$_0$ and in the sense of Weihrauch
reducibility, to the following version of the Rainbow Ramsey
Theorem, which was shown by Patey \cite{P} to be strictly weaker than
TS$^2$.

\begin{defn}
RRT$^2_2$: Let $c : [\mathbb N]^2 \rightarrow \mathbb N$ be such that
$|c^{-1}(i)| \leq 2$ for all $i$. Then there is an infinite set $R$
such that $c$ is injective on $[R]^2$.
\end{defn}

As discussed in \cite{CDHJSW}, the proof of Theorem \ref{rsthm}
carries through in RCA$_0$, from which it will follow that so does the
proof of Lemma \ref{mainlem} that we will give below. Thus the proof
of Theorem \ref{dncthm} also carries through in RCA$_0$, except for
one issue: Having $|W_e^{\emptyset'}| \geq m$ does not necessarily
imply in RCA$_0$ that $E^m_e$ is defined. (The issue is that RCA$_0$
does not imply the $\Pi^0_1$-bounding principle.) However, we can get
around this problem exactly as in Section 4 of \cite{CDHJSW}, by using
the principle 2-EI defined there, thus obtaining the following.

\begin{thm}
\textup{thin-HT}$^{=2}$ implies \textup{RRT}$^2_2$ over \textup{RCA}$_0$. 
\end{thm}

We can also obtain a Weihrauch reduction from RRT$^2_2$ to a
version of thin-HT$^{=2}$ as in the final paragraph of Section 4 of
\cite{CDHJSW}, but we have to be a bit careful because in the proof of
Theorem \ref{dncthm}, the function witnessing that $S$ is effectively
immune relative to $\emptyset'$ is obtained uniformly not from $S$,
but from an $n$ such that $c(x+y) \neq n$ for all distinct $x,y \in
S$. Let strong thin-HT$^{=2}$ be the version of thin-HT$^{=2}$
where a solution to an instance $c$ consists of both a solution $S$ to
$c$ as an instance of thin-HT$^{=2}$ and an $n$ as above. Then we have
the following.

\begin{thm}
\textup{RRT}$^2_2$ is Weihrauch-reducible to \textup{strong
thin-HT}$^{=2}$.
\end{thm}

We do not know, however, whether this theorem remains true if we
replace strong thin-HT$^{=2}$ by thin-HT$^{=2}$.

None of the above results depend on the addition function in
particular, and can be adapted as in \cite{CDHJSW} to any function $f:
[\mathbb N]^2 \rightarrow \mathbb N$ that is \emph{addition-like},
which means that
\begin{enumerate}

\item $f$ is computable,

\item there is a computable function $g$ such that $f(\{x,y\})>n$ for
all $y>g(x,n)$, and

\item there is a $b$ such that for all $x \neq y$, there are at most
$b$ many $z$'s for which $f(\{x,z\})=f(\{x,y\})$.

\end{enumerate}

We finish this section by proving Lemma \ref{mainlem}.

\begin{proof}[Proof of Lemma \ref{mainlem}]
Let $f$ be a binary function and $D$ an $f$-large set. We will apply
Theorem \ref{rsthm} to obtain a computable $c : \mathbb N \rightarrow
2$. We then define $D^i = \{n \in D : c(n)=i\}$. The value of $q$ will
not matter here, so let us fix $q=\frac{1}{2}$. Let $M$ be as in
Theorem \ref{rsthm}.

Let $g$ be a computable injective binary function with computable
image such that $kg(e,k) \leq 2^{\frac{g(e,k)}{2}}$ and $g(e,k) \geq
M$ for all $e$ and $k$.

Say that $s$ is \emph{acceptable for $e,k$} if $|D \cap (s +
E^{f(e,kg(e,k))}_e[s])| \geq kg(e,k)$ and for every $t<s$ such that $(s
+ E^{f(e,kg(e,k))}_e[s]) \cap (t + E^{f(e,kg(e,k))}_e[t]) \neq
\emptyset$, we have $ E^{f(e,kg(e,k))}_e[s] = E^{f(e,kg(e,k))}_e[t]$. If
$s$ is acceptable for $e,k$ then let $F_{e,k,s,0}$ be the first
$g(e,k)$ many elements of $s + E^{f(e,kg(e,k))}[s]$, let $F_{e,k,s,1}$ be
the next $g(e,k)$ many elements of $s + E^{f(e,kg(e,k))}[s]$, and so on,
until $F_{e,k,s,k-1}$.

Let $\mathcal F$ consist of all $F_{e,k,s,j}$ for all $e,k$, all $s$
acceptable for $e,k$, and all $j<k$. Then we can arrange the elements
of $\mathcal F$ into a computable sequence of finite sets, each of
size at least $M$. Fix $x$ and $m$. If $m$ is not in the image of
$g$ then there are no elements of $\mathcal F$ of size $m$. Otherwise,
there is a unique pair $e,k$ such that $m=g(e,k)$, and all elements of
$\mathcal F$ of size $m$ that contain $x$ are of the form
$F_{e,k,s,j}$ for some $s \leq x$. We can computably determine all
such sets from $m$ and $x$, and the definition of acceptability means
that there are at most $kg(e,k) \leq 2^{\frac{m}{2}}$ many such sets.

Thus the hypotheses of Theorem \ref{rsthm} hold, and hence there is a
$c$, obtained uniformly computably from $f$ and $D$, such that none of
the sets in $\mathcal F$ are homogeneous for $c$. Let
$\widehat{f}(e,k)=f(e,kg(e,k))$ and let $D^i = \{n \in D :
c(n)=i\}$. Fix $e$ and $k$ such that $E_e^{\widehat{f}(e,k)}$ is
defined. If $s$ is sufficiently large then $s$ is acceptable for
$e,k$, and $F_{e,k,s,j} \subseteq s + E_e^{\widehat{f}(e,k)}$ for all
$j<k$. For each $j<k$ and $i<2$, there is at least one $x \in
F_{e,k,s,j}$ such that $c(x)=i$. Since the $F_{e,k,s,j}$ are disjoint,
$|D^i \cap (s + E_e^{\widehat{f}(e,k)})| \geq k$. Thus $D^i$ is
$\widehat{f}$-large.
\end{proof}

\section{Open Questions}
\label{open}

In this section, we collect a few open questions and possible
directions for further work arising from the above results.

\begin{q}
Does thin-HT imply ACA$_0$ over RCA$_0$ (i.e., without assuming
I$\Sigma^0_2$)?
\end{q}

Of course, one way to give a positive answer to this question would be
to show that thin-HT implies I$\Sigma^0_2$ over RCA$_0$. If that is
not the case, then it could be interesting to try to determine the
first-order part of thin-HT.

\begin{q}
Is thin-HT provable in ACA$_0$?
\end{q}

\begin{q}
Does thin-HT imply HT, say over RCA$_0$?
\end{q}

In the spirit of Hindman, Leader, and Strauss~\cite{HLS}, we can also
ask the less formal question of whether there is a proof of thin-HT
that is not already a proof of HT.

\begin{q}
Is RRT$^2_2$ Weihrauch-reducible to thin-HT$^{=2}$ (as opposed to
strong thin-HT$^{=2}$)?
\end{q}

\begin{q}
What is the exact relationship between thin-HT$^{=2}$ and each of
TS$^2$, RRT$^2_2$, and HT$^{=2}$?
\end{q}

There are also versions of the Thin Set Theorem for colorings with
finitely many colors. For example, an instance of TS$^n_k$ is a
coloring $c$ of $[\mathbb N]^n$ with $k$ many colors, and a solution
to this instance is an infinite set $T$ such that $|c([T]^n)|<k$. This
principle and RT$^n_k$ form the two ends of a spectrum of principles
RT$^n_{k,j}$ for $1 \leq j < k$, where an instance is a coloring $c$
of $[\mathbb N]^n$ with $k$ many colors, and a solution to this
instance is an infinite set $T$ such that $|c([T]^n)| \leq j$. It
would be interesting to pursue versions of HT based on these
principles. One might hope to show, for instance, that there is a
boundary between principles that ``behave like HT'', e.g.\ HT$^{\leq
  2}_4$, which as mentioned in the introduction was shown to imply
ACA$_0$ in \cite{CKLZ}; and those that ``behave like versions of TS /
RT'', e.g.\ the thin version of HT$^{\leq 2}_4$, which can easily be
shown to follow from RT$^2_{4,2}$.

\end{document}